OpenStudy (anonymous):

One valve drains a pool in 2 hours, and another drains it in 9 hours. How long will it take to drain the pool with both valves open at the same time? Round to the nearest tenth. 4.5 hours 7 hours 1.6 hours 0.6 hours

OpenStudy (ganpat):

let A drain in 2 hrs.. So A drains in unit time = 1/2 and B drains in unit time = 1/9 so together they will drain in A + B = 1/2 + 1/9 = (9 + 2 ) / 18 = 18/11 = 1.63 = 1.6 hrs...

OpenStudy (ganpat):

u understood ??

OpenStudy (anonymous):

yeah i got it

OpenStudy (ganpat):

great..